Project

General

Profile

Actions

Commissioning and Ops Scheduling meeting : 9/2/2016

To be held in CR10 in the ATLAS Building @ 15:30 pm

Phone access : ReadyTalk number
US 866-740-1260
UK 0800-496-0576
Swiss 0800-70-53-51
Italy 800182592
Access Code: 2155402

Agenda:

  • Review of tasks and activities identified at last meeting
    • Action Item : John will continue to investigate pathologies in scattering plots
    • Action Item : Pierrick/Mark - ensure that the pressure and material in the absorber volumes are recorded
    • Action Item : Steve to talk to Alan Stevens about stopping 4-hourly monitoring but leaving the alarm in place
      • Alan replied and is quite happy to stop the 4 hourly checks, but keep the alarm in the ISIS CR
    • Action Item : Colin, Trevor, Steve, Pierrick to have a meeting to review the Daresbury work schedule in light of running time
      and the currently known state of the SS work. Meeting to happen on morning of Thursday 28th.
      • Action stands
    • Action Item : Steve/DC/Andy to think about organisation during mixed work sessions
      • Action stands
    • Action Item Steve/Colin to talk to the tracker group and request action plan
      • Tracker group contacted and report will be made at the meeting. See discussion below
    • Action Item Andy/Colin/DC to think about the procedure of running with SC permit with the PRY open (@ 10 A and with
      sanction from the US team)
      • A procedure exists. We should check that it has been run past Andy.
  • Current and near future activities
    • DS Cooldown and PS return schedule
    • Preparation for realignment work
    • Focus coil installation and vacuum test schedule
    • Tracker patch panel leak work and schedule
    • Controls and Monitoring : Network reconfiguration and testing
    • Daresbury work schedule
  • List of things to address before February run
  • Operational Schedule
    • Expert led run : Feb 17th. Hall to be left closed overnight
    • Beam tuning from 10 am on Feb 18th
    • Data-taking nominally begins on the 23rd with shift training on the 22nd.
  • AOB
  • Chair of meetings for next 5 weeks

Minutes

Present: Josef B, Paul H, Vicky B, Steve B, Paul K, P. Franchini, Mark T, John C, Ed O, John G,
Henry N, Pierrick H, Ryan B

  • Review of tasks and activities identified at last meeting
    • Action Item : John will continue to investigate pathologies in scattering plots
      • John is looking into this but no result at this meeting
    • Action Item : Pierrick/Mark - ensure that the pressure and material in the absorber volumes are recorded
      • A pull-down list on run control will let the shifters define the absorber to be used in each run
      • pressure sensors are available and will be read out into CDB
    • Action Item : Steve to talk to Alan Stevens about stopping 4-hourly monitoring but leaving the alarm in place
      • Alan replied and is quite happy to stop the 4 hourly checks, but keep the alarm in the ISIS CR
    • Action Item : Colin, Trevor, Steve, Pierrick to have a meeting to review the Daresbury work schedule in light of running time
      and the currently known state of the SS work. Meeting to happen on morning of Thursday 28th.
      • Action stands
    • Action Item : Steve/DC/Andy to think about organisation during mixed work sessions
      • Action stands
    • Action Item Steve/Colin to talk to the tracker group and request action plan
      • see below
    • Action Item Andy/Colin/DC to think about the procedure of running with SC permit with the PRY open (@ 10 A and with
      sanction from the US team)
      • Action stands
  • Current and near future activities
    • DS Cooldown and PS return schedule
      • Decay solenoid is cooling down with no problems. It is expected to be cold around the end of this week.
      • DS PS will return by end of next week. Trevor would like Mike to inspect the PS before installation on the mezzanine.
      • Commissioning is planned for 29th Feb - 3rd Mar, and 8th-11th Mar
      • Calibration of current will be done during tests
    • Preparation for realignment work
      • MIPO have agreed to postpone the manual realignment until the end of the user cycle
    • Focus coil installation and vacuum test schedule
      • Jon Duffy working on FC1 in the Hall and getting it ready and he is close. This will
        take until the end of this week.
      • FC can't be put into the channel until the work on the tracker leak is done however.
    • Tracker patch panel leak work and schedule
      • (Long long discussion, summarised as)
      • There are three areas where He leaks from the tracker : the Hall probe connectors, the patch panel and the diffuser.
      • Leaks from the connectors and patch panel will be dealt with this week
      • Repair of diffuser leak is more involved and requires expert effort from Oxford. We need quantitative estimates of the
      • leak rate to judge how important this is.
      • Tenatively plan diffuser leak repair in the day of the 28th Feb - 3rd Mar.
    • Controls and Monitoring : Network reconfiguration and testing
      • Pierrick is rebuilding all the IOCs. We will finish this week so will be able test by 16th.
      • Intention is to be ready for the expert led run on the 17th.
    • Daresbury work schedule
      • See above
      • SS dump trays to be taken back to Daresbury this week
  • Operational Schedule
    • Expert led run : Feb 17th. Hall to be left closed overnight
    • Beam tuning from 10 am on Feb 18th
    • Data-taking nominally begins on the 23rd with shift training on the 22nd.
  • AOB
    • Henry reminded us that the Hall walkthrough on the 22nd should also include RF crew
    • Paul informed us that there will be a visit by school children sometime in March and will need a Hall tour.
  • Chair of meetings for next 5 weeks
    • Steve is teaching at this time for the next 5 weeks. DC will chair the meeting.

Updated by Boyd, Steven over 7 years ago ยท 18 revisions